OpenStudy (nathanielborrero321):

The formula for the volume of a sphere is V = 4/3 ∏r3, where r is the radius of the sphere. The radius of the Moon is 1/4 of the radius of the Earth. Answer the following questions. a. If the volume of the Earth is V = 4/3 ∏r^3, write the volume of the Moon in terms of r. b. Find the ratio of the volume of the Earth to the volume of the Moon. Can someone help me with this question?

OpenStudy (math&ing001):

The radius of the Moon is 1/4 of the radius of the Earth If the volume of the Earth is V = 4/3 ∏r^3 then r is the radius of the Earth. So radius of the moon is 1/4 r Plug that in the formula to get the volume of the moon.

OpenStudy (nathanielborrero321):

Ok let me do that

OpenStudy (nathanielborrero321):

1.45?

OpenStudy (math&ing001):

You can't possibly get a value by plugging (1/4 r) into the formula. You should get a volume in terms of r. Here's a start : \[V=\frac{ 4 }{ 3 }\pi r^3=\frac{ 4 }{ 3 }\pi (\frac{1}{4} r)^3\] Then simplify.

OpenStudy (nathanielborrero321):

ok

OpenStudy (nathanielborrero321):

so then it would be v= 0.07r^3?

OpenStudy (math&ing001):

It's better if you keep it in fraction form.

OpenStudy (nathanielborrero321):

ok

OpenStudy (math&ing001):

To find the ratio Earth:Moon you put \[\frac{volume of earth}{volume of moon}\]

OpenStudy (nathanielborrero321):

I got 64.71

OpenStudy (math&ing001):

For ratio ?

OpenStudy (nathanielborrero321):

4.53:0.07 Is the ratio

OpenStudy (nathanielborrero321):

I'm really confused

OpenStudy (math&ing001):

What's 64.71 for ?

OpenStudy (nathanielborrero321):

I divided 4.53 by 0.07 for some reason. Its probably irrelevant.

OpenStudy (math&ing001):

OK, let's get back to where we were. What did you get for the volume of the moon in fraction form

OpenStudy (nathanielborrero321):

4/3*\[\pi\]1/64 r^3

OpenStudy (math&ing001):

It doesn't show very well for me, can you rewrite ?

OpenStudy (nathanielborrero321):

yea

OpenStudy (waterlover9150802):

4/3 * pi* 1/64r^3

OpenStudy (nathanielborrero321):

oh thanks water

OpenStudy (waterlover9150802):

np

OpenStudy (math&ing001):

Yes but still needs simplification (simplify 4 and 64)

OpenStudy (waterlover9150802):

1/48 *pi*r^3

OpenStudy (nathanielborrero321):

I also got that

OpenStudy (waterlover9150802):

oh ok

OpenStudy (math&ing001):

Yep, correct ! Now for the last part \[ratio=\frac{ volume~~of~~the~~earth }{ volume~~of~~the~~moon }\]

OpenStudy (nathanielborrero321):

YAY

OpenStudy (nathanielborrero321):

4/3*pi*r^3:1/48 *pi*r^3

OpenStudy (nathanielborrero321):

then simplify it to be...

OpenStudy (nathanielborrero321):

64/1

OpenStudy (math&ing001):

Yes ! 64:1 is the ratio !

OpenStudy (nathanielborrero321):

Sorry for taking so much of your time

OpenStudy (math&ing001):

Not.At.All ! It's a pleasure =)